Important Notes


                   
1. All the plotted data and images are just for browsing purpose and 
   are not fully calibrated.  Please do not use the data at any 
   presentations and publications without contacting the Principal 
   Investigators, K. Shiokawa (shiokawa at stelab.nagoya-u.ac.jp).

2. Digital data are available on request.  Please contact the Principal 
   Investigator to copy the data.  The Principal Investigator
   may limit the use of the data in order to avoid possible confliction of 
   the scientific directions of the users.   

3. The all-sky image data at Rikubetsu (camera 3 after October, 1998) is 
   upside down (upward is toward the south) for October 20, 1998 - March 31,
   1999.  Please see figure captions.  

4. For the SATI data, the values of temperature may be corrected in the 
   future by using new calibration data sets.  Please also see the background 
   level in the plots.  High background means cloudy sky during which the 
   temperature and intensity data become meaningless.  

5. Please refer  Shiokawa et al. [1999] for details of the instrumentation.  



**************** some additional local notes *******

1. Camera A (10) at Paratunka was stopped for March 5-11, 2008, 
   due to condensation in the lens system after severe cyclone 
   on February 28, 2008.  The lens was cleaned on August 16-17, 2008. 

2. The CCD temperature of FP01 at Tromso raised from -80C (normal) 
   to ~-65C at Feb.16-Mar.11, 2010 due to water flow degradation.
   So the 630-nm temperature data may be affected. 

3. The FPI04 at Darwin has scanning problem.  The sky scanner does 
   not move for May 21-July 7, 2011, due to missing the position 
   sensor.  

4. The Camera #11 (Camera B) at Magadan has a problem of A/D converter 
   of CCD since early 2014.  We make fast-scan measurements since April 2015. 
   The sensitivity becomes ~1/16.5 from slow scan to fast scan, and 
   the satulation level becomes from 65535 counts to 4096 counts.  

5. The Camera #10 (camera A) at Paratunka was repaired due to CCD cooler
   trouble and restarted the observation from 16 June 2015. 



OMTI Home Page .


Division II Home Page .


STELAB Home Page .